OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation x - 4y = 4. slope = 4; y intercept at (0, -1) slope = ; y intercept at (0, 1) slope = ; y intercept at (0, -1) slope = -4; y intercept at (0, 1)

OpenStudy (gorv):

4y=x-4 y=x/4-1 slope=1 y intercept put x=0 y=-1

OpenStudy (anonymous):

Thanks gorv

OpenStudy (gorv):

@manueld167 welcome buddy

OpenStudy (anonymous):

You need to set it up in y = mx + b form, where m is always your slope and b is your y intercept (when x is 0). Start by subtracting the x. -4y = -x + 4 Divide the -4 over y = (-x +4)/-4 simplifies to y = x/4 - 1 compare that to y - mx + b b = -1 (you y intercept (0,-1) and m = 1/4 (your slope). I type slower, but I hope the detail helps you understand it better. My apologies